OpenStudy (anonymous):

Help! if a store offers 3 for the price of 2. What is the implied discount? a) 20% b) 33% c) 40% d) 66%

jimthompson5910 (jim_thompson5910):

how far did you get?

OpenStudy (anonymous):

i'm really bad a doing percentage's. i have no idea where to start @jim_thompson5910

jimthompson5910 (jim_thompson5910):

let's say each item is $20 how much would it cost to buy 2 items? 3 items?

OpenStudy (anonymous):

for 2 it would be $40 and for 3 it would be $60.

jimthompson5910 (jim_thompson5910):

how much do you save if you buy 3 items but you pay the price as if it were 2 items

OpenStudy (anonymous):

i would save $20. is it 20%?

jimthompson5910 (jim_thompson5910):

you do save $20 but it's not 20%

jimthompson5910 (jim_thompson5910):

it's only 20% IF the total price for the three items was $100

jimthompson5910 (jim_thompson5910):

you save $20. This is the discount divide the discount by the original price to get 20/60 = 1/3 = 0.3333... = 33.33%

jimthompson5910 (jim_thompson5910):

the discount as a fixed number is $20 the discount as a percentage (of the original price) is roughly 33% the idea is that you take 33% of the original price ($60) and subtract that off the original price

OpenStudy (anonymous):

ok, now i understand how to do this. thank you so much, again!

jimthompson5910 (jim_thompson5910):

no problem